# Moiré Superradiance in Cavity Quantum Electrodynamics with Quantum Atom Gas

### v2

@@ VI. Atom diffusion dynamics
$\frac{d\alpha}{dt}$ $=\left[i\left(\Delta_{c}-u\operatorname{Re}\left[\psi^{\ast}\mathcal{M}_{2f_{n+1}}\psi\right]\right)-\kappa\right]\alpha-i\eta\operatorname{Re}\left[\psi^{\ast}\mathcal{M}_{f_{n+1}}\psi\right]+\xi/\sqrt{N},$
$\frac{d\psi}{dt}$ $=-i\left[\omega_{k}\mathcal{K}+u\left|\alpha\right|^{2}\left(2\mathds{1}+\mathcal{M}_{2f_{n+1}}\right)+\frac{V_{l}}{4}\mathcal{M}_{2f_{n}}+\eta\operatorname{Re}(\alpha)\mathcal{M}_{f_{n+1}}\right]\psi,$ (S17)
in which $\left\{{\alpha(0),\ \psi(0)}\right\}$ are sampled from $\mathcal{W}(\alpha,\alpha^{\ast},\psi,\psi^{\ast},0)$ and $\xi$ is a complex Gaussian random number with variance $\kappa$ and mean value 00. Specifically, we sample with
$\left(\begin{array}[c]{c}\alpha\\ \psi_{i}\end{array}\right)=\left(\begin{array}[c]{c}\left(m+in\right)/2\sqrt{N}\\ o_{i}+\left(p_{i}+iq_{i}\right)/2\sqrt{N}\end{array}\right),$ (S18)
@@
VI atom diffusion dynamics
We have adopted two different schemes to study atomic diffusion and the physics discovered by both schemes are in qualitative agreement. In both schemes the cavity field is treated using TWA with the details described in V. Difference lies in that in one scheme the atomic wavefunction is simulated with Gross-Pitaevskii equation
$i\frac{d\Psi\left(x,t\right)}{dt}=\left[-\frac{\hbar^{2}}{2m}\partial_{x}^{2}+4u\left|\alpha\right|^{2}\cos^{2}\left(k_{c}x\right)+V_{l}\cos^{2}\left(k_{l}x\right)+2\eta\operatorname{Re}(\alpha)\cos\left(k_{c}x\right)\right]\Psi\left(x,t\right),$ (S19)
@@
$\phi\left(x\right)=\left(w\sqrt{2\pi}\right)^{-1/2}\exp\left[-\frac{\left(x-L/2\right)^{2}}{4w^{2}}\right],$ (S20)
@@
$\phi\left(x\right)=a_{0}\sqrt{\frac{1}{L}}+\underset{N\rightarrow\infty}{\lim}\sqrt{\frac{2}{L}}\left({\displaystyle\sum\limits_{n=1}^{N}}a_{n}\cos\left(2\pi nx\right)+{\displaystyle\sum\limits_{n=1}^{N}}b_{n}\sin\left(2\pi nx\right)\right),$ (S21)
@@
$a_{0}=\sqrt{\sqrt{2\pi}2w}\operatorname{erf}\left(\frac{1}{4w}\right),a_{n}=2\sqrt{\sqrt{2\pi}w}\cos\left(n\pi\right)e^{-\left(2n\pi w\right)^{2}}\operatorname{Re}\left[\operatorname{erf}\left(\frac{1}{4w}+2in\pi w\right)\right],$ (S22)
@@
with $\operatorname{erf}(\cdot)$ the error function. Compare (S21) with the mode expansion (3) in the main text, in TWA sampling (S18) we choose $o_{i}=a_{i}$.
@@
Figure S1: Original atomic wavepacket (left column) compared with those from reconstruction (right column). The top row is for $w=0.02$, while the second row is for $w=0.1$.
@@
Note that as we have omitted the sine term and truncated at $n_{c}$, these will affect the atomic wavepacket reconstructed after TWA simulation, especially when the wavepacket width is small. This is indicated in Fig. S1, in which we compared the original wavepacket with the reconstructed one. However this only cause quantitative differences and will not affect our main results in the text.
Figure S2: Quantum diffusion dynamics calculated with TWA. (a) Atom wavepacket width at steady state $w_{s}$ versus pumping strength $\eta$. In the shaded region $w_{s}$ at $M=3$ is larger than that at $M=5$, with the insets demonstrating localized (extended) atomic momentum distribution for $M=3$ ($5$) respectively at a $\eta$ marked by the black dashed line. The corresponding evolution dynamics of atomic wavepacket width $w$ and cavity photon number are given in (b). The parameters are the same as those in the main text.
@@
With periodic boundary condition considered here, as shown in Fig. S2(b), the width increases at initial time and then gradually approaches a steady value. In the meanwhile, the cavity photon number jump from zero before reaching saturation. We plot the steady state width $w_{s}$ versus the pumping strength in Fig. S2(a). As expected, since the lattice constant becomes larger with increased $M$, one will have broader wavepacket in the $M=5$ case as compared with $M=3$ for an initial identical wavepacket. This is true at small $\eta$. $w_{s}$ abruptly decreases upon the pumping strength passing across a critical value. This diffusion behavior tuned between two distinct diffusive regimes arises from superradiant phase transitions: (i) For small pumping strengths, cavity field is not excited and atomic wavepacket spreads until it reach the unit cell boundary, with the width showcasing damped oscillation as illustrated by the green line in Fig. S2(b). In the steady state, atomic wavefunction distribution is localized in momentum space as shown in the upright inset of Fig. S2(a), indicating extended distribution in coordinate space. (ii) While at large pumping, intracavity field build up to coform a moiré lattice. Different atomic momentum modes are populated via scattering, resulting in momentum space extended distribution at steady state (downleft inset of Fig. S2(a)), and henceforce suppressed expansion and width.
@@
Noteworthy that in the shaded region of Fig. S2(a), for the very same pumping strength, one counter-intuitively have larger $w_{s}$ in the $M=3$ case instead of $M=5$. In this specific parameter region, atomic wavepacket diffuse in distinct different manner determined by the moiré parameter: In the $M=3$ case atomic diffusion behavior can be depicted by (i), while the atomic diffusion in the $M=5$ case is attributed to (ii). This completely opposite behavior is manipulated solely by a change in $M$ with all the other physical parameters identical, thus can be regarded as moiré effect. One can understand this phenomenon as the result of interplay between quantum interference (Anderson localization) and quantum fluctuations of the cavity field. The aid of cavity is indispensable to observe the moiré effect.

### v3

@@ Appendix E Truncated Wigner methods
$\frac{d\alpha}{dt}$ $=\left[i\left(\Delta_{c}-u\operatorname{Re}\left[\psi^{\ast}\mathcal{M}_{2f_{n+1}}\psi\right]\right)-\kappa\right]\alpha$
$-i\eta\operatorname{Re}\left[\psi^{\ast}\mathcal{M}_{f_{n+1}}\psi\right]+\xi/\sqrt{N},$
$\frac{d\psi}{dt}$ $=-i\left[\omega_{k}\mathcal{K}+u\left|\alpha\right|^{2}\left(2\mathds{1}+\mathcal{M}_{2f_{n+1}}\right)+\frac{V_{l}}{4}\mathcal{M}_{2f_{n}}\right$.
$\left.+\eta\operatorname{Re}(\alpha)\mathcal{M}_{f_{n+1}}\right]\psi,$ (37)
in which $\left\{{\alpha(0),\ \psi(0)}\right\}$ are sampled from $\mathcal{W}(\alpha,\alpha^{\ast},\psi,\psi^{\ast},0)$ and $\xi$ is a complex Gaussian random number with variance $\kappa$ and mean value 0. Specifically, we sample with
$\left(\begin{array}[]{c}\alpha\\ \psi_{i}\end{array}\right)=\left(\begin{array}[]{c}\left(m+in\right)/2\sqrt{N}\\ o_{i}+\left(p_{i}+iq_{i}\right)/2\sqrt{N}\end{array}\right),$ (38)
@@
Appendix F Atom diffusion dynamics
We have adopted two different schemes to study atomic diffusion and the physics discovered by both schemes are in qualitative agreement. In both schemes the cavity field is treated using TWA with the details described in E. Difference lies in that in one scheme the atomic wavefunction is simulated with Gross-Pitaevskii equation
$i\frac{d\Psi\left(x,t\right)}{dt}$ $=\left[-\frac{\hbar^{2}}{2m}\partial_{x}^{2}+4u\left|\alpha\right|^{2}\cos^{2}\left(k_{c}x\right)+V_{l}\cos^{2}\left(k_{l}x\right)\right$.
$\left.+2\eta\operatorname{Re}(\alpha)\cos\left(k_{c}x\right)\right]\Psi\left(x,t\right),$ (39)
@@ Appendix F Atom diffusion dynamics
$\phi\left(x\right)=\left(w\sqrt{2\pi}\right)^{-1/2}\exp\left[-\frac{\left(x-L/2\right)^{2}}{4w^{2}}\right],$ (40)
@@
$\phi\left(x\right)$ $=a_{0}\sqrt{\frac{1}{L}}+\underset{N\rightarrow\infty}{\lim}\sqrt{\frac{2}{L}}\left(\sum\limits_{n=1}^{N}a_{n}\cos\left(2\pi nx\right)\right$.
$\left.+\sum\limits_{n=1}^{N}b_{n}\sin\left(2\pi nx\right)\right),$ (41)
@@
$a_{0}$ $=\sqrt{\sqrt{2\pi}2w}\operatorname{erf}\left(\frac{1}{4w}\right),$
[…]
$\left.\left.+2in\pi w\right)\right],$ (42)
@@
with $\operatorname{erf}(\cdot)$ the error function. Compare (41) with the mode expansion (3) in the main text, in TWA sampling (38) we choose $o_{i}=a_{i}$.
@@
Figure 6: Original atomic wavepacket (left column) compared with those from reconstruction (right column). The top row is for $w=0.02$, while the second row is for $w=0.1$.
@@
Note that as we have omitted the sine term and truncated at $n_{c}$, these will affect the atomic wavepacket reconstructed after TWA simulation, especially when the wavepacket width is small. This is indicated in Fig. 6, in which we compared the original wavepacket with the reconstructed one. However this only cause quantitative differences and will not affect our main results in the text.
@@
With periodic boundary condition considered here, as shown in Fig. 7(b), the width increases at initial time and then gradually approaches a steady value. In the meanwhile, the cavity photon number jump from zero before reaching saturation. We plot the steady state width $w_{s}$ versus the pumping strength in Fig. 7(a). As expected, since the lattice constant becomes larger with increased $M$, one will have broader wavepacket in the $M=5$ case as compared with $M=3$ for an initial identical wavepacket. This is true at small $\eta$. $w_{s}$ abruptly decreases upon the pumping strength passing across a critical value. This diffusion behavior tuned between two distinct diffusive regimes arises from superradiant phase transitions: (i) For small pumping strengths, cavity field is not excited and atomic wavepacket spreads until it reach the unit cell boundary, with the width showcasing damped oscillation as illustrated by the green line in Fig. 7(b). In the steady state, atomic wavefunction distribution is localized in momentum space as shown in the upright inset of Fig. 7(a), indicating extended distribution in coordinate space. (ii) While at large pumping, intracavity field build up to coform a moiré lattice. Different atomic momentum modes are populated via scattering, resulting in momentum space extended distribution at steady state (downleft inset of Fig. 7(a)), and henceforce suppressed expansion and width.
Figure 7: Quantum diffusion dynamics calculated with TWA. (a) Atom wavepacket width at steady state $w_{s}$ versus pumping strength $\eta$. In the shaded region $w_{s}$ at $M=3$ is larger than that at $M=5$, with the insets demonstrating localized (extended) atomic momentum distribution for $M=3$ (55) respectively at a $\eta$ marked by the black dashed line. The corresponding evolution dynamics of atomic wavepacket width $w$ and cavity photon number are given in (b). The parameters are the same as those in the main text.
@@
Noteworthy that in the shaded region of Fig. 7(a), for the very same pumping strength, one counter-intuitively have larger $w_{s}$ in the $M=3$ case instead of $M=5$. In this specific parameter region, atomic wavepacket diffuse in distinct different manner determined by the moiré parameter: In the $M=3$ case atomic diffusion behavior can be depicted by (i), while the atomic diffusion in the $M=5$ case is attributed to (ii). This completely opposite behavior is manipulated solely by a change in $M$ with all the other physical parameters identical, thus can be regarded as moiré effect. One can understand this phenomenon as the result of interplay between quantum interference (Anderson localization) and quantum fluctuations of the cavity field. The aid of cavity is indispensable to observe the moiré effect.